# Perturbation-based Analysis of Compositional Data

### v1

@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
We can extend the considerations of Example 3 to more general amalgamating perturbations. For disjoint $A,B\subseteq[d]$, consider an $(A\to B)$-amalgamating perturbation. Suppose as before that $Z:=\mathsf{C}(X)$ for some $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale. We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation so we suppose now that $A\cup B=[d]$. The multiplicative operation described in Example 3 can be generalized by letting $u_{B}=\sum_{j\in B}e_{j}$ and considering
@@
$\phi=(\phi^{L},\phi^{W}):z\mapsto(-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}})$
@@ 2.1.3 Perturbations arising from alternative simplex geometries
When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$. While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions. We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature. The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison (1982) who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data. A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ by
@@
This transformation maps the interior of $\Delta^{d-1}$ to the subspace of $\mathbb{R}^{d}$ given by $\{x\in\mathbb{R}^{d}\mid\sum_{j=1}^{d}x^{j}=0\}$. The inverse CLR is defined for all $x\in\operatorname{Im}(\mathrm{CLR})$ by
@@
which can be normalized to a unit $1$-norm vector. The direction of this derivative can be vastly different from the centering direction used in Example 5. To obtain a derivative-isolating reparametrization, we can instead define a unit-speed perturbation on the log-ratio scale, $\widetilde{\psi}:(x,\gamma)\mapsto x-\gamma\frac{x}{\|x\|_{2}}$, and find a reparametrization on this scale. It can be shown that on the log-ratio scale, a derivative-isolating reparametrization is given by $\phi^{L}:x\mapsto-\|x\|_{2}$ and $\phi^{W}(z):=x/\|x\|_{2}$ and we can transform these to reparametrizations of $Z$ by composing with the $\mathrm{CLR}$-transform.
@@
$\partial_{\gamma}\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j})^{k}\bigm{|}_{\gamma=0}=\mathbbm{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
@@ 2.2.1 Effects of changes in individual components
As a first example we consider perturbation effects that capture how a response is associated with changes in individual components. These types of effects are, for example, useful as a variable selection tool in compositional data. In Section 4.3, we illustrate this for a human gut microbiome study. For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured. We then wants to determine for which microbes a high abundance is associated with a high BMI. The simplex constraint implies that for all coordinates $j$ the conditional independence
@@
is trivially satisfied and feature importance or influence measures based on (11) are meaningless. Intuitively, this stems from the fact that changing a single component always requires a change in at least some of the remaining components. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. To that end, we already considered these perturbations in Example 3 where we proposed two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed (see Example 3 for more intuition on these speeds). $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$. A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
@@ 2.2.2 Effects of changes in diversity
As a second example we consider perturbation effects that capture how a response is associated with increases in diversity. Such effects are often relevant in social sciences and econometrics, e.g., when considering racial or income distributions. As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response. Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (the centering perturbations defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$). We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 5 such that $L$ is equal to the negative Gini coefficient. The construction in Example 5 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics. In Section 4.2, we use data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression.
@@ 2.2.3 Effects of amalgamations
As a third example we consider perturbation effects that capture the influence of general amalgamations. As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness). Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness. As above, we are again interested in effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$. However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally. For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic. To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports. Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$. We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short. As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example 4).
@@ 2.3 Perturbations as causal quantities
Perturbations are a priori non-causal quantities and should not be confused with (causal) interventions. To make this distinction clear, we now discuss how our perturbation framework can be used for causal inference. Causal models (Pearl, 2009; Rubin, 2005) provide a mathematical framework for modeling changes to the data generating mechanism – often called interventions. In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data generating mechanism under hypothetical interventions. The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution. Unlike interventions, perturbations describe hypothetical changes to the predictor space (in this work $\Delta^{d-1}$, but the same ideas hold more generally) and not to the data-generating mechanism. To make this more precise and show how to use perturbations for causal inference, we use the structural causal model (SCM) framework (Pearl, 2009; Bongers et al., 2021). Similar considerations, however, also apply with the potential outcome framework.
@@
Assume now we observe data generated by $\mathcal{M}$ and are interested in understanding how $Z$ causally affects $Y$. One target parameter might be to consider the expected value of $Y$ under an intervention that sets $Z$ to a fixed value, using do-notation (Pearl, 2009), this can be expressed as $\mathbb{E}[Y\mid\operatorname{do}(Z=z)]=\mathbb{E}[g(z,H)]$. For larger $d$ the full causal relation might be too ambitious of a goal, and it may be more tractable to look at the effect of individual components. Since the vector $Z$ is compositional, one needs to be careful when expressing summary interventions. Perturbations provide a solution to this as they allow us to express sensible ways of changing the predictors. More specifically, let $\psi$ be a perturbation as defined in Definition 1. Then, for all $\gamma\in[0,1]$, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$. In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$. Here, we only consider interventions on the entire compositional vector $Z$, which means that it is not possible to model how causal effects propagate within the composition itself. The corresponding causal versions of the target parameters in Definition 2 are then the causal average binary perturbation effect
@@
Unlike their non-causal versions, the causal average perturbation effects may not be identifiable from the observational distribution. Therefore, to estimate them, we need to make additional causal assumptions. For example, we may assume the variable $(H,Y,Z)$ were generated by an SCM with the same causal structure as $\mathcal{M}$. In that case, if $H$ is observed, the expectation of $Y$ under perturbation induced interventions on $Z$ can be expressed as
@@
and hence is identifiable. The formula in (12) is a special case of the adjustment formula (e.g., Pearl, 2009), but other ways of identifying this effect also exist (e.g., propensity score matching). It is easy to adapt the estimators proposed in this work to such settings by conditioning on additional confounding covariates $H$. In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$. This is for example done in the numerical experiment in Section 4.2.
@@ 3 Estimating perturbation effects
In this section, we describe how to estimate and test hypotheses about the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$. We will not explicitly restrict ourselves to any one perturbation $\psi$ as the approaches described are agnostic to this choice. However, we will assume that any directional perturbation comes with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and for any binary perturbation we define $\phi^{L}:z\mapsto\mathbbm{1}_{\{}z=\psi(z,1)\}$ and $\phi^{W}:z\mapsto\psi(z,1)$. This allows us to transform our estimation problem from a compositional data problem where we observe $n$ i.i.d. copies of $(Y,Z)\in\mathbb{R}\times\Delta^{d-1}$ to a generic problem with $n$ observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$. We only present an overview of the different estimators in this section and postpone a detailed description of the algorithms and theory to Section 5.
@@ 3.1 Efficient estimation using semiparametric theory
There are two main issues with this estimate; (a) repeated usage of the same data for both the estimation of $f$ and the computation of the target (double-dipping) and (b) estimation bias inherited from the bias in $\widehat{f}$. Issue (a) can be resolved by splitting the dataset into two parts of roughly equal size and using the first half to compute $\widehat{f}$ and the second part to compute $\widetilde{\tau}_{\psi}$. To increase finite-sample performance of this splitting procedure, one can swap the roles of the two halves and average the resulting estimates. This procedure is known as cross-fitting. Sometimes it is preferable to split in more than two folds to increase the sample size in the estimation of $\widehat{f}$ and to derandomize the resulting estimate. Issue (b) is more complicated to resolve and requires the machinery of semiparametric theory. One well-established approach is to start from the plug-in estimator and derive a bias-correction, known as a one-step correction, that can be added to the plug-in estimator and weakens the conditions under which the estimator is asymptotically Gaussian. The one-step correction can be found by computing the influence function of the target parameter, which can be thought of as analogous to a derivative when the target parameter is viewed as a function defined on a space of distributions. For more details on this general procedure, see e.g. Kennedy (2023). It can be shown that these bias-corrected estimators are asymptotically efficient, meaning that, at least asymptotically, they have the smallest possible variance of any estimator. Influence functions have already been computed for the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$ that we estimate here and we will describe the details of these estimates in the following sections. Figure 3 contains an illustration of the coverage of confidence intervals resulting from target parameter and variance estimates from the different methods mentioned above.

### v2

@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
We can extend the considerations of Example 3 to more general amalgamating perturbations. For disjoint $A,B\subseteq[d]$, consider an $(A\to B)$-amalgamating perturbation. Suppose as before that $Z:=\mathsf{C}(X)$ for some $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale. We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation, so we suppose now that $A\cup B=[d]$. The multiplicative operation described in Example 3 can be generalized by letting $u_{B}=\sum_{j\in B}e_{j}$ and considering
@@
$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}\right)$
@@ 2.1.3 Perturbations arising from alternative simplex geometries
When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$. While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions. We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature. The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison (1982) who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data. A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ (the interior of the simplex) by
@@
This transformation maps $(\Delta^{d-1})^{\circ}$ to the subspace of $\mathbb{R}^{d}$ given by $\{x\in\mathbb{R}^{d}\mid\sum_{j=1}^{d}x^{j}=0\}$. The inverse CLR is defined for all $x\in\operatorname{Im}(\mathrm{CLR})$ by
@@
which can be normalized to a unit $1$-norm vector. The direction of this derivative can be vastly different from the centering direction used in Example 5. To obtain a derivative-isolating reparametrization, we can instead define a unit-speed perturbation on the log-ratio scale (normalizing with the Euclidean norm as is natural here), $\widetilde{\psi}:(x,\gamma)\mapsto x-\gamma\frac{x}{\|x\|_{2}}$, and find a reparametrization on this scale. It can be shown that on the log-ratio scale, a derivative-isolating reparametrization is given by $\phi^{L}:x\mapsto-\|x\|_{2}$ and $\phi^{W}(z):=x/\|x\|_{2}$ and we can transform these to reparametrizations of $Z$ by composing with the $\mathrm{CLR}$-transform.
@@
$\partial_{\gamma}\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j})^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
@@ 2.2.1 Effects of changes in individual components
As a first example we consider perturbation effects that capture how a response is associated with changes in individual components. These types of effects are, for example, useful as a variable selection tool in compositional data. In Section 4.3, we illustrate this for a human gut microbiome study. For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured. We then want to determine for which microbes a high abundance is associated with a high BMI. The simplex constraint implies that for all coordinates $j$ the conditional independence
@@
is trivially satisfied and feature importance or influence measures based on (11) are meaningless. Intuitively, this stems from the fact that changing a single component always requires a change in at least some of the remaining components. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. To that end, we already considered these perturbations in Example 3 where we proposed two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed (see Example 3 for more intuition on these speeds). $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S6 of the supplementary material). A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
@@ 2.2.2 Effects of changes in diversity
As a second example we consider perturbation effects that capture how a response is associated with increases in diversity. Such effects are often relevant in social sciences and econometrics, e.g., when considering racial or income distributions. As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response. Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (the centering perturbations defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$). We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 5 such that $L$ is equal to the negative Gini coefficient. The construction in Example 5 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics. In Section 4.2.1, we create a semi-synthetic dataset using data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression. In Section 4.2.2 we analyze the effect of racial diversity on grades in New York schools to further illustrate the use of $\mathrm{CDI}_{\mathrm{Gini}}$.
@@ 2.2.3 Effects of amalgamations
As a third example we consider perturbation effects that capture the influence of general amalgamations. As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness). Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness. As above, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$. However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally. For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic. To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports. Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$. We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short. As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example 4).
@@ 2.3 Perturbations as causal quantities
Perturbations are a priori non-causal quantities and should not be confused with (causal) interventions. To make this distinction clear, we now discuss how our perturbation framework can be used for causal inference. Causal models (Pearl, 2009; Rubin, 2005) provide a mathematical framework for modeling changes to the data-generating mechanism – often called interventions. In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data-generating mechanism under hypothetical interventions. The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution. Unlike interventions, perturbations describe hypothetical changes to the predictor space (in this work $\Delta^{d-1}$, but the same ideas hold more generally) and not to the data-generating mechanism. To make this more precise and show how to use perturbations for causal inference, we use the structural causal model (SCM) framework (Pearl, 2009; Bongers et al., 2021). Similar considerations, however, also apply with the potential outcome framework.
@@
Assume now we observe data generated by $\mathcal{M}$ and are interested in understanding how $Z$ causally affects $Y$. One target parameter might be to consider the expected value of $Y$ under an intervention that sets $Z$ to a fixed value, using do-notation (Pearl, 2009), this can be expressed as $\mathbb{E}[Y\mid\operatorname{do}(Z=z)]=\mathbb{E}[g(z,H)]$. For larger $d$ the full causal relation might be too ambitious of a goal, and it may be more tractable to look at the effect of individual components. Since the vector $Z$ is compositional, one needs to be careful when expressing summary interventions. Perturbations provide a solution to this as they allow us to express sensible ways of changing the predictors. More specifically, let $\psi$ be a perturbation as defined in Definition 1. Then, for all $\gamma$ sufficiently small, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$. In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$. Here, we only consider interventions on the entire compositional vector $Z$, which means that it is not possible to model how causal effects propagate within the composition itself. The corresponding causal versions of the target parameters in Definition 2 are then the causal average binary perturbation effect
@@
Unlike their non-causal versions, the causal average perturbation effects may not be identifiable from the observational distribution. Therefore, to estimate them, we need to make additional causal assumptions. For example, we may assume the variables $(H,Y,Z)$ were generated by an SCM with the same causal structure as $\mathcal{M}$. In that case, if $H$ is observed, the expectation of $Y$ under perturbation induced interventions on $Z$ can be expressed as
@@
and hence is identifiable. The formula in (12) is a special case of the adjustment formula (e.g., Pearl, 2009), but other ways of identifying this effect also exist (e.g., propensity score matching). It is easy to adapt the estimators proposed in this work to such settings by conditioning on additional confounding covariates $H$. In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$. This is for example done in the numerical experiments in Section 4.2.
@@ 3 Estimating perturbation effects
In this section, we describe how to estimate and test hypotheses about the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$. We will not explicitly restrict ourselves to any one perturbation $\psi$ as the approaches described are agnostic to this choice. However, we will assume that any directional perturbation comes with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and for any binary perturbation we define $\phi^{L}:z\mapsto\mathds{1}_{\{z=\psi(z,1)\}}$ and $\phi^{W}:z\mapsto\psi(z,1)$. This allows us to reparametrize our estimation problem from a compositional data problem with $n$ i.i.d. observations of $(Y,Z)\in\mathbb{R}\times\Delta^{d-1}$ into a generic problem with $n$ i.i.d. observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$. We only present an overview of the different estimators in this section and postpone a detailed description of the algorithms and theory to Section 5.
@@ 3.1 Efficient estimation using semiparametric theory
There are two main issues with this estimate; (a) repeated usage of the same data for both the estimation of $f$ and the computation of the target (double-dipping) and (b) estimation bias inherited from the bias in $\widehat{f}$. Issue (a) can be resolved by splitting the dataset into two parts of roughly equal size and using the first half to compute $\widehat{f}$ and the second part to compute $\widetilde{\tau}_{\psi}$. To increase finite-sample performance of this splitting procedure, one can swap the roles of the two halves and average the resulting estimates. This procedure is known as cross-fitting. Sometimes it is preferable to split in more than two folds to increase the sample size in the estimation of $\widehat{f}$ and to derandomize the resulting estimate. Issue (b) is more complicated to resolve and requires the machinery of semiparametric theory. One well-established approach is to start from the plug-in estimator and derive a bias-correction, known as a one-step correction, that can be added to the plug-in estimator and weakens the conditions under which the estimator is asymptotically Gaussian. The one-step correction can be found by computing the influence function of the target parameter, which can be thought of as analogous to a derivative when the target parameter is viewed as a function defined on a space of distributions. For more details on this general procedure, see e.g. Kennedy (2023). It can be shown that these bias-corrected estimators are asymptotically efficient, meaning that, at least asymptotically, they have the smallest possible variance of any estimator. Influence functions have already been computed for the reparametrized target parameters $\tau_{\psi}$ and $\lambda_{\psi}$ that we estimate here and we will describe the details of these estimates in the following sections. Figure 3 contains an illustration of the coverage of confidence intervals resulting from target parameter and variance estimates from the different methods mentioned above.